# Convergent Privacy Framework with Contractive GNN Layers for Multi-hop Aggregations

### v1

@@ 2.2. Differential Privacy for GNNs
Intuitively, DP (Dwork et al., 2006) theoretically quantifies the privacy of a model by measuring the indistinguishability of the outputs of a mechanism $\mathcal{M}$ on two adjacent datasets $D$ and $D^{\prime}$. It can be classified into bounded DP and unbounded DP depending on the construction of $D^{\prime}$, where the former is by replacing a data instance of $D$ and the latter is by addition / removal of a data sample of $D$. The privacy budget $\epsilon$ is smaller representing a stronger privacy guarantee, while $\delta$ is a slackness quantity that relaxes the pure DP constraint.
2.2.1. Privacy Definition on Graphs
@@
Definition 0 (Edge-level adjacency (Karwa et al., 2011)).
@@
Definition 0 (Node-level adjacency (Karwa et al., 2011)).
@@
2.2.2. Perturbed Message Passing with DP
To incorporate differential privacy into GNNs, one can add noise to the message passing layer, following the perturbed message passing approach (Sajadmanesh et al., 2023). Given a graph ${\bm{G}}$ and message passing function $\mathrm{MP}$, we define a sequence $\{{\bm{X}}^{(k)}\}_{k=0}^{K}$ of node feature matrices by:
(2) ${\bm{X}}^{(k+1)}=\Pi_{\mathcal{K}}(\mathrm{MP}_{{\bm{G}}}({\bm{X}}^{(k)})+{\bm{Z}}^{(k)})$
@@ 2.2.2. Perturbed Message Passing with DP
(3) $\Delta(\mathrm{MP})=\max_{G,G^{\prime}}\max_{{\bm{X}}\in\mathcal{K}}\|\mathrm{MP}_{{\bm{G}}}({\bm{X}})-\mathrm{MP}_{{\bm{G}}^{\prime}}({\bm{X}})\|_{F}$
@@
Privacy Accounting.
Privacy accounting process analyzes the total privacy budget for the composition of several (adaptive) private algorithms. A common approach for analyzing the Gaussian mechanism in perturbed message passing is through Rényi differential privacy (RDP) (Mironov, 2017) and its composition theorem.
Definition 0 (Rényi differential privacy (Mironov, 2017)).
A randomized algorithm $\mathcal{M}$ is $(\alpha,\epsilon)$-RDP for $\alpha>1$, $\epsilon>0$ if for every adjacent datasets $X,X^{\prime}$, we have $D_{\alpha}(\mathcal{M}(X)\|\mathcal{M}(X^{\prime}))\leq\epsilon$, where $D_{\alpha}(P\|Q)$ is the Rényi divergence of order $\alpha$ between probability distributions $P$ and $Q$ defined as:
@@
Theorem 6 (Composition of RDP (Mironov, 2017)).
@@
In this work, we present our privacy results in terms of RDP for ease of interpretation, while our underlying analysis employs the tighter $f$-DP framework. This analysis leverages recent advances in privacy amplification techniques (Schuchardt et al., 2024; Bok et al., 2024) to achieve stronger privacy guarantees. The technical details of our convergent privacy analysis are discussed in §3.2.
3. Multi-hop Private GNNs Initiative
Deeper GNNs (Thürlemann and Riniker, 2023; Wu et al., 2021a) are vital in tasks like modeling molecular structures, where some properties depend on long range interactions (Dwivedi et al., 2022; Chen et al., 2024b; Tönshoff et al., 2024) of the nodes (Zhang et al., 2022c; Li et al., 2023b). In this long-range scenario, message passing GNNs must flow across multiple hops before reaching a target node—requiring stacking multiple layers to exchange information across $K$-hop neighborhoods. However, ensuring privacy in multi-layer GNNs poses key challenges. In this section, we outline and illustrate these challenges via a motivating case study (Section 3.1), ultimately motivating our new design insights (Section 3.2) based on contractive message passing.
3.1. Observations on Privacy Accumulation and Performance Degradation
A Case Study on GAP’s Performance on Learning Long-Range Interactions.
To further investigate this phenomenon, we evaluate GAP on the chain-structured dataset (see Figure 2), which is adopted in Gu et al. (2020); Liu et al. (2021) to examine long-range interaction learning capabilities. Specifically, this dataset creates a controlled environment for evaluating GNN performance on long-range interactions, as an MLP model would necessarily fail since most nodes have zero-valued feature vectors regardless of their chain type. Consequently, for a message-passing GNN to correctly classify nodes positioned $K$ hops away from the informative first node, it must perform at least $K$ propagation steps to transfer the meaningful features across the chain as shown in Figure 2. This requirement becomes particularly challenging in the private setting as noise must be injected after each message passing layer, potentially overwhelming the signal being propagated.
Figure 2. Message Passing on Chain-structured Dataset
Figure 3 compares GAP’s non-private accuracy (blue solid line) with its private versions (dashed lines), under different privacy budgets $\epsilon\in\{1,2,4,8,16,32\}$. It shows a binary node classification over multiple $8$-node chains as a case study. The red solid line in Figure 3 represents random guessing (50% accuracy).
Figure 3. Motivating Experiments of Classification Model over Chain-structured Dataset.
@@ 3.1. Observations on Privacy Accumulation and Performance Degradation
Figure 4. Overview of $\mathsf{CARIBOU}$. The framework integrates three key modules: the Contractive Aggregation Module (CAM) for generating embeddings with adjustable coefficients, the Privacy Allocation Module (PAM) for efficient privacy budget allocation, and the Privacy Auditing Module (PDM) for empirical privacy verification. Combining CAM and PAM can form the Perturbed Contractive Message Passing (PCMP). Together, these modules enable scalable and accurate private GNN learning under a convergent privacy budget.
Empirical (Counter-Intuitive) Observation: Deeper GNNs May Enhance Privacy. Recent work by Wu et al. (2024) revealed a counter-intuitive phenomenon: deeper GNNs empirically exhibit lower vulnerability to membership and link inference attacks. This challenges standard privacy composition analysis, which suggests privacy risk increases with model depth due to multiple queries of the graph data. The insight stems from the over-smoothing [^1] Incidentally, another similar phrase is “over-fitting”, which refers to a model performing well on training data but poorly on unseen data due to memorization. We emphasize that they are different concepts to avoid conflation. (Chen et al., 2020a) phenomenon, where node representations become indistinguishably homogeneous phenomenon, where node representations converge toward similar values as depth increases, making it inherently difficult for adversaries to distinguish individual nodes or infer sensitive relationships.
@@
3.2. Our Approach: Contractive Message Passing for Convergent Privacy
Our key insight is to identify and leverage the inherent privacy amplification that occurs in deeper GNNs — a phenomenon that appears counter-intuitive but can be formally explained through the mathematical properties of message passing operations. This privacy amplification is intrinsically connected to the well-known over-smoothing phenomenon in GNN. Over-smoothing (Chen et al., 2020a) emerges fundamentally from the contractiveness of typical graph aggregation operations. When nodes aggregate information from their neighbors (e.g., through averaging as in GCN), the resulting representations necessarily become more similar to each other. This mathematical property can be formally defined:
@@ 3.2. Our Approach: Contractive Message Passing for Convergent Privacy
[…]
Designing Private GNNs with Convergent Privacy.
We aim to leverage the privacy amplification properties of contrative map to design a new framework for private GNNs. This framework is inspired by the recent advances in differentially private gradient descent (DP-GD) (Chourasia et al., 2021; Bok et al., 2024), which has shown that the privacy cost can converge to a finite value even with arbitrarily many iterations. Motivated by it, we aim to translate the advanced privacy analysis techniques from DP-GD to GNNs. We observe that perturbed message passing (Equation 2) in GNNs follows a strikingly mathematical parallel pattern as DP-GD: This parallel structure enables to leverage the insight of convergent privacy analysis in the context of GNNs, provided we ensure two critical conditions:
  1. (1)
    Hidden intermediate embeddings: Release only the final node representations ${\bm{X}}^{(K)}$ after $K$ layers, concealing all intermediate states; (Section 4.1)
  2. (2)
    Contractive message passing: Design the message passing operation $\mathrm{MP}_{{\bm{G}}}$ to be provably contractive with coefficient $c<1$, ensuring $\|\mathrm{MP}_{{\bm{G}}}({\bm{X}})-\mathrm{MP}_{{\bm{G}}}({\bm{Y}})\|_{F}\leq c\|{\bm{X}}-{\bm{Y}}\|_{F}$ for all node feature matrices ${\bm{X}},{\bm{Y}}$. (Section 4.2)
@@
4. Private GNNs with Contractiveness
Building on insights in Section 3, we aim to design a message-passing layer that is contractive with a flexibility to adjust the Lipschitz constant, enabling a more efficient and effective privacy budget allocation. To this end, we propose $\mathsf{CARIBOU}$, a modular private GNN framework (Figure 4) with three key modules:
@@ 4. Private GNNs with Contractiveness
    Contractive Aggregation Module (CAM, § 4.2): We propose a new design of message passing layers with carefully controlled Lipschitz constants to ensure contractiveness so that we can add a reasonable, small yet sufficient noise to protect the computed message passing; Then, we realize a new DP mechanism of perturbed message passing that only releases final node representations ${\bm{X}}^{(K)}$, preventing adversary from exploiting intermediate states and thus amplifying the privacy guarantee rooted in hidden node embeddings.
@@
    Privacy Allocation Module (PAM, § 4.3): PAM ensures efficient privacy budget allocation for edge- and node-level DP guarantees, in which noise calibration is based on new convergent privacy analysis.
@@
    Privacy Auditing Module (PDM, § 4.4): PDM empirically audits graph DP through tests such as membership inference attacks, ensuring theoretical guarantees align with practical deployments.
By integrating these modules, $\mathsf{CARIBOU}$ (Figure 4) achieves accurate and scalable private GNN learning on long-range graphs, with adjustable edge- and node-level privacy under a convergent privacy budget which we will theoretically established in Section 5. Table 6 summarizes acronyms and symbols.
4.1. Clarification on Security Model
Typically, aggregation-based GNNs like GCN (Kipf and Welling, 2017), GCN-II (Chen et al., 2020b), and SAGE (Hamilton et al., 2017) adopt finally obtained node embeddings, while GAP (Sajadmanesh et al., 2023) utilizes the concatenation of all intermediate obtained node embeddings of $K$ layers. This work based on message passing mechanisms adopts the former route to generate node embeddings. Notably, the implicit (also realistic) assumption (Ye and Shokri, 2022) here is that the internal state of $K$-layer message passing is not revealed to the adversary when we design DP algorithms. The final node representation can be applied later in various tasks, including node classification (Wang et al., 2024), link prediction (Rao et al., 2024), or graph classification (Li et al., 2024). Current privacy analysis in private GNNs (e.g., GAP) implies the assumption of revealing all $K$-layer node embeddings computed by perturbed message passing mechanisms to the adversary, which grows the noise scale $\sigma$ linearly ($O(\sqrt{{K}/{\epsilon^{2}}})$) for deriving a theoretical privacy bound. Such privacy bound is loose and we refine it under a more realistic security model (Ye and Shokri, 2022).
@@ 4.1. Clarification on Security Model
Table 1. Acronyms & Symbols
4.2. Contractive Aggregation Module
Contractive operations in GNNs, such as the graph convolution (GConv) layer (Kipf and Welling, [n. d.]), inherently reduce privacy risks by mitigating the memorization of GNN parameters through the over-smoothing phenomenon (Chen et al., 2020a). This property aligns well with the need for $K$-layer aggregation in long-range graphs, where a target node aggregates embeddings from distant source nodes. However, directly stacking $K$ GConv layers, as shown in Figure 3, is suboptimal due to limited expressive power (Chen et al., 2020b) and heightened sensitivity to DP noise (Wu et al., 2019).
@@ 4.2. Contractive Aggregation Module
(4) ${\bm{X}}^{(k+1)}=C_{\mathsf{L}}\big{(}\alpha_{1}\hat{{\bm{A}}}{\bm{X}}^{(k)}+\alpha_{2}\mathsf{Mean}({\bm{X}}^{(k)})\big{)}+\beta{\bm{X}}^{(0)},$
@@
Analysis of hyperparameters.
The hyperparameter $C_{\mathsf{L}}$ ensures the contractiveness by bounding the magnitude of the output embeddings. Specifically, $C_{\mathsf{L}}$ enforces a Lipschitz constraint, ensuring that small changes in the input embeddings ${\bm{X}}^{(k)}$ result in proportionally small changes in the output ${\bm{X}}^{(k+1)}$. The coefficients $\alpha_{1},\alpha_{2}$ govern the relative contributions of the graph-based aggregation $\hat{{\bm{A}}}{\bm{X}}^{(k)}$ and the mean-based aggregation $\mathsf{Mean}({\bm{X}}^{(k)})$, respectively. By satisfying the constraint $\alpha_{1}+\alpha_{2}=1$, these coefficients ensure a convex combination of the two components. A higher $\alpha_{1}$ emphasizes the influence of the graph topology, leveraging structural information from the adjacency matrix $\hat{{\bm{A}}}$. The parameter $\beta$ controls the residual connection $\beta{\bm{X}}^{(0)}$ calculated independently of the graph topology, incorporating the characteristics of the initial node ${\bm{X}}^{(0)}$ into the output. This residual term mitigates the vanishing gradient problem by preserving a direct path for gradient flow.
@@
Comparison with existing on general design.
GAP (Sajadmanesh et al., 2023) enhances expressiveness by connecting all intermediate embeddings but lacks contractiveness, compromising utility and privacy. In contrast, CGL introduces adjustable coefficients, achieving a balanced trade-off between privacy, utility, and generalizability. Table 2 summarizes the general design characteristics between GAP and $\mathsf{CARIBOU}$.
4.3. Privacy Allocation Module

### v2

@@ II-B Differential Privacy for GNNs
Intuitively, DP [tcc/DworkMNS06] theoretically quantifies the privacy of a model by measuring the indistinguishability of the outputs of a mechanism $\mathcal{M}$ on two adjacent datasets $D$ and $D^{\prime}$. It can be classified into bounded DP and unbounded DP depending on the construction of $D^{\prime}$, where the former is by replacing a data instance of $D$ and the latter is by addition / removal of a data sample of $D$. The privacy budget $\epsilon$ is smaller representing a stronger privacy guarantee, while $\delta$ is a slackness quantity that relaxes the pure DP constraint.
II-B1 Privacy Definition on Graphs
@@
Definition 2 (Edge-level adjacency [pvldb/KarwaRSY11]).
@@
Definition 3 (Node-level adjacency [pvldb/KarwaRSY11]).
@@
II-B2 Perturbed Message Passing with DP
To incorporate differential privacy into GNNs, one can add noise to the message passing layer, following the perturbed message passing approach [uss/sajadmanesh2023gap]. Given a graph $\bm{G}$} and message passing function $\mathrm{MP}$, we define a sequence $\{{\bm{X}}^{(k)}\}_{k=0}^{K}$ of node feature matrices by:
${\bm{X}}^{(k+1)}=\Pi_{\mathcal{K}}(\mathrm{MP}_{{\bm{G}}}({\bm{X}}^{(k)})+{\bm{Z}}^{(k)})$ (2)
@@ II-B2 Perturbed Message Passing with DP
$\Delta(\mathrm{MP})=\max_{G,G^{\prime}}\max_{{\bm{X}}\in\mathcal{K}}\|\mathrm{MP}_{{\bm{G}}}({\bm{X}})-\mathrm{MP}_{{\bm{G}}^{\prime}}({\bm{X}})\|_{F}$ (3)
@@
II-B3 Privacy Accounting
Privacy accounting process analyzes the total privacy budget for the composition of several (adaptive) private algorithms. A common approach for analyzing the Gaussian mechanism in perturbed message passing is through Rényi differential privacy (RDP) [csf/mironov2017renyi] and its composition theorem.
Definition 5 (Rényi differential privacy [csf/mironov2017renyi]).
A randomized algorithm $\mathcal{M}$ is $(\alpha,\epsilon)$-RDP for $\alpha>1$, $\epsilon>0$ if for every adjacent dataset $X,X^{\prime}$, we have $D_{\alpha}(\mathcal{M}(X)\|\mathcal{M}(X^{\prime}))\leq\epsilon$, where $D_{\alpha}(P\|Q)$ is the Rényi divergence of order $\alpha$ between probability distributions $P$ and $Q$ defined as:
@@ II-B3 Privacy Accounting
Theorem 1 (Composition of RDP [csf/mironov2017renyi]).
@@
In this work, we present our privacy results in terms of RDP for ease of interpretation, while our underlying analysis employs the tighter $f$-DP framework. This analysis leverages recent advances in privacy amplification techniques [nips/SchuchardtSKG24, icml/BokShifted24] to achieve stronger privacy guarantees. The technical details of our convergent privacy analysis are discussed in §III-B.
Figure 2: Message Passing on Chain-structured Dataset
III Private Multi-layer GNNs Initiative
Multi-layer GNNs [iclr/ThurlemannR23, nips/WuJWMGS21] are vital in tasks like modeling molecular structures, where some properties depend on long range interactions [nips/DwivediRGPWLB22, corr/abs-2406-03386, tmlr/TonshoffRRG24] of the nodes [zhang2022deep, li2023neural]. Specifically, they require messages flowing across multiple hops before reaching a target node through stacking multiple layers to exchange information across $K$-hop neighborhoods. However, ensuring privacy in multi-layer GNNs poses key challenges. In this section, we outline and illustrate these challenges via a motivating case study (Section III-A), ultimately motivating our new design insights (Section III-B) based on contractive message passing.
III-A Observations on Privacy Accumulation and Performance Degradation
III-A1 A Case Study on GAP’s Performance on Learning Long-Range Interactions
To further investigate this phenomenon, we evaluate $K$-layer GAP on the chain-structured dataset (see Figure 2), which is adopted in [nips/GuC0SG20, liu2021eignn] to examine long-range interaction learning capabilities. Specifically, this dataset creates a controlled environment for evaluating GNN performance on long-range interactions. A model without learning from graph structure such as MLP would fail since most nodes in the dataset have zero-valued feature vectors regardless of their chain type. For a message-passing GNN to correctly classify nodes positioned $K$ hops away from the informative first node, it must perform at least $K$ propagation steps to transfer the meaningful features across the chain as shown in Figure 2. This requirement becomes particularly challenging in the private setting as noise must be injected after each message passing layer, potentially overwhelming the signal being propagated.
Figure 3 compares GAP’ accuracy of its non-private version (blue solid line) with its private versions (dashed lines), under different privacy budgets $\epsilon\in\{1,2,4,8,16,32\}$. It shows a binary node classification over multiple $8$-node chains as a case study. The red solid line in Figure 3 represents random guessing (50% accuracy).
Figure 3: Motivating Experiments of Classification Model over Chain-structured Datasets.
@@ III-A1 A Case Study on GAP’s Performance on Learning Long-Range Interactions
Figure 4: Overview of $\mathsf{CARIBOU}$. The framework integrates CAM, PAM, PDM (Section IV). Combining CAM and PAM can form PCMP. Together, these modules enable scalable and accurate private GNN learning under a convergent privacy budget.
III-A2 Empirical (Counter-Intuitive) Observation: Deeper GNNs May Enhance Privacy
Recent work by [wu2024provable] revealed a counter-intuitive phenomenon: deeper GNNs empirically exhibit lower vulnerability to membership and link inference attacks. This challenges standard privacy composition analysis, which suggests privacy risk increases with model depth due to multiple queries of the graph data. The insight stems from the over-smoothing [^1] Incidentally, another similar phrase is “over-fitting”, which refers to a model performing well on training data but poorly on unseen data due to memorization. We emphasize that they are different concepts to avoid confusion. [aaai/ChenLLLZS20] phenomenon, where node representations become indistinguishably homogeneous phenomenon, where node representations converge toward similar values as depth increases, making it inherently difficult for adversaries to distinguish individual nodes or infer sensitive relationships.
@@ III-A2 Empirical (Counter-Intuitive) Observation: Deeper GNNs May Enhance Privacy
III-B Core Idea for Convergent Privacy
Our insight is to identify and leverage the inherent privacy amplification that occurs in multi-layer GNNs through contractiveness (Definition 6). When nodes aggregate information from their neighbors (e.g., graph convolution), the resulting representations necessarily become more similar to each other.
@@ III-B Core Idea for Convergent Privacy
A map $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is said to be contractive with respect to a norm $\|\cdot\|$ if there exists a constant $c<1$ such that for all $x,y\in\mathbb{R}^{d}$: $\|f(x)-f(y)\|\leq c\|x-y\|$ where $c$ is the contractiveness coefficient that governs the rate of contraction.
Designing Private GNNs with Convergent Privacy. We aim to leverage the privacy amplification properties of contractive map to design a new framework for private GNNs. This framework is inspired by the recent advances in differentially private gradient descent (DP-GD) [nips/ChourasiaYS21, icml/BokShifted24], which has shown that the privacy cost can converge to a finite value even with arbitrarily many iterations. Motivated by it, we aim to translate the advanced privacy analysis techniques from DP-GD to GNNs. We observe that perturbed message passing (Equation 2) in GNNs follows a strikingly mathematical parallel pattern as DP-GD: This parallel structure enables to leverage the insight of convergent privacy analysis in the context of GNNs, provided we ensure two critical conditions:
  1. 1.
    Hidden intermediate embeddings: Release only the final node representations ${\bm{X}}^{(K)}$ after $K$ layers, concealing all intermediate states; (Section IV)
  2. 2.
    Contractive message passing: Design the message passing operation $\mathrm{MP}_{{\bm{G}}}$ to be provably contractive with coefficient $c<1$, ensuring $\|\mathrm{MP}_{{\bm{G}}}({\bm{X}})-\mathrm{MP}_{{\bm{G}}}({\bm{Y}})\|_{F}\leq c\|{\bm{X}}-{\bm{Y}}\|_{F}$ for all node feature matrices ${\bm{X}},{\bm{Y}}$. (Section IV-A)
@@
IV Private GNNs with Contractiveness
Building on insights in Section III, we propose $\mathsf{CARIBOU}$, a modular private GNN framework with convergent privacy budget (Figure 4), including three key modules:
@@ IV Private GNNs with Contractiveness
    Contractive Aggregation Module (CAM, § IV-A): We propose a new design of message passing layers with carefully controlled Lipschitz constants to ensure contractiveness so that we can add a reasonable, small yet sufficient noise to protect the computed message passing; Then, we realize a new DP mechanism of perturbed message passing that only releases final node representations ${\bm{X}}^{(K)}$, preventing adversary from exploiting intermediate states and thus amplifying the privacy guarantee rooted in hidden node embeddings.
@@
    Privacy Allocation Module (PAM, § IV-B): PAM ensures efficient privacy budget allocation for edge- and node-level DP guarantees, in which noise calibration is based on new convergent privacy analysis.
@@
    Privacy Auditing Module (PDM, § IV-C): PDM empirically audits graph DP through tests such as membership inference attacks, ensuring theoretical guarantees align with practical deployments.
Security Model. Aggregation-based GNNs such as GCN [iclr/kipf2017semi], GCN-II [icml/ChenWHDL20], and SAGE [nips/HamiltonYL17] reveal only the final node embeddings, keeping intermediate embeddings hidden. The final node embeddings learned in this DP manner can be subsequently applied to downstream tasks such as node classification [www/WangLDLW24], link prediction [www/RaoKR24], and graph classification [icdm/LiW024]. Since DP introduces utility loss, GAP [uss/sajadmanesh2023gap] seeks to improve model utility by concatenating all intermediate node embeddings from the $K$ layers. However, it causes the noise scale $\sigma$ to grow linearly as $O(\sqrt{{K}/{\epsilon^{2}}})$, and this privacy bound is loose. This work takes a step further beyond GAP, removing the assumption that all intermediate node embeddings must be revealed under a more realistic security model [nips/0001S22].
Overall, $\mathsf{CARIBOU}$ (Figure 4) achieves accurate and private multi-layer GNNs, which theory will be established in Section V. Table VI summarizes acronyms and symbols.
@@
TABLE II: Acronyms & Symbols
IV-A Contractive Aggregation Module
Contractive operations in GNNs, such as the graph convolution (GConv) layer [iclr/KipfW17], inherently reduce privacy risks by mitigating the memorization of GNN parameters through the over-smoothing phenomenon [aaai/ChenLLLZS20]. This property aligns well with the need for $K$-layer aggregation in long-range graphs, where a target node aggregates embeddings from distant source nodes. However, directly stacking $K$ GConv layers, as shown in Figure 3, is suboptimal due to limited expressive power [icml/ChenWHDL20] and heightened sensitivity to DP noise [WuSZFYW19].
@@ IV-A Contractive Aggregation Module
${\bm{X}}^{(k+1)}=C_{\mathsf{L}}\big(\alpha_{1}\hat{{\bm{A}}}{\bm{X}}^{(k)}+\alpha_{2}\mathsf{Mean}({\bm{X}}^{(k)})\big)+\beta{\bm{X}}^{(0)},$ (4)
@@
Analysis of hyperparameters. The hyperparameter $C_{\mathsf{L}}$ ensures the contractiveness by bounding the magnitude of the output embeddings. Specifically, $C_{\mathsf{L}}$ enforces a Lipschitz constraint, ensuring that small changes in the input embeddings ${\bm{X}}^{(k)}$ result in proportionally small changes in the output ${\bm{X}}^{(k+1)}$. The coefficients $\alpha_{1},\alpha_{2}$ govern the relative contributions of the graph-based aggregation $\hat{{\bm{A}}}{\bm{X}}^{(k)}$ and the mean-based aggregation $\mathsf{Mean}({\bm{X}}^{(k)})$, respectively. By satisfying the constraint $\alpha_{1}+\alpha_{2}=1$, these coefficients ensure a convex combination of the two components. A higher $\alpha_{1}$ emphasizes the influence of the graph topology, leveraging structural information from the adjacency matrix $\hat{{\bm{A}}}$. The parameter $\beta$ controls the residual connection $\beta{\bm{X}}^{(0)}$ calculated independently of the graph topology, incorporating the characteristics of the initial node ${\bm{X}}^{(0)}$ into the output. This residual term mitigates the vanishing gradient problem by preserving a direct path for gradient flow.
@@
Comparison with existing design. GAP [uss/sajadmanesh2023gap] enhances expressiveness by connecting all intermediate embeddings but lacks contractiveness, compromising utility and privacy. In contrast, CGL introduces adjustable coefficients, achieving a balanced trade-off between privacy, utility, and generalizability. Table I summarizes the general design characteristics between GAP and $\mathsf{CARIBOU}$.
IV-B Privacy Allocation Module